Accounting - Sven Enterprises

Seven Enterprises is a large producer of gourmet pet food. During April, it produced 147 batches of puppy meal. Each batch weighs 1,000 pounds. To produce thisquantity of output, the company purchased and used 148,450 pounds of direct materials at a cost of $593,800. IT ALSO INCURRED DIRECT LABOR COST OF $17, 600 for the2,200 hours worked by employees on the puppy meal crew. Manufacturing overhead incurred at the puppy meal plant during April totaled $3,625, of which $2,450 wasconsidered fixed. Seven’s standard cost information for 1,000-pound batches of puppy meal is as follows:

direct materials standard price………………………. $4.20 per pound
standard quantity allowed per batch……………… 1,020 pounds
direct labor standard rate………………………………… $8.50 per hour
Standard hours allowed per batch…………………… 14 direct labor hours
Fixed overhead budgeted………………………………… $2,800 per month
Normal level of production…………………………. 140 batches per month
Variable overhead application rate…………… $9.00 per batch
Fixed overhead application rate
($2,800/140 batches)………………………………… 20.00 per batch
Total overhead application rate………………… $29.00 per batch

Instructions
a. compute the materials price and quantity variances
b. Compute the labor rate and efficiency variances
c. Compute the manufacturing overhead spending and volume variance
d. Record the journal entry to charge materials (at standard) to work in process
e. Record the journal entry to charge direct labor (at standard) to work in process.
f. Record the journal entry to charge manufacturing overhead ( at standard) to work in process
g. Record the journal entry to transfer the 147 batches of puppy meal produced in April to finished goods.
h. Record the journal entry to close any over- or under applied overhead to cost of goods sold
Answer #1

Materials Price Variance= Actual Quantity Used × (Standard Price − Actual Price)

= 148450( 4.20 – 4) = 29690 Favorable

Actual Price per Pound = 593800 / 148450 = 4 pounds

Materials Quantity Variance = Standard Price × (Standard Quantity − Actual Quantity)

= 4.20 (149,940 pounds -148450) =6258 Favorable

Standard Quantity Allowed = 147 batches × 1,020 pounds/batch = 149,940 pounds

Labor Rate Variance= Actual Labor Hours × (Standard Rate − Actual Rate)

2200 x(8.50 -8) = 1100 Favorable

Actual Rate per Hour = 17600 / 2200 =$8.00/hour

Labor Efficiency Variance = Standard Hourly Rate × (Standard Hours − Actual Hours)

= 8.50 x(2058- 2220)= -$1,207 (or $1,207 Unfavorable)

Standard Hours Allowed = 1 47x 14= 2058 hours

answered by: Gudda